# Meta-Learning Guided Pruning for Few-Shot Plant Pathology on Edge Devices

### v1

@@ Abstract
Farmers in remote areas need quick and reliable methods for identifying plant diseases, yet they often lack access to laboratories or high-performance computing resources. Deep learning models can detect diseases from leaf images with high accuracy, but these models are typically too large and computationally expensive to run on low-cost edge devices such as Raspberry Pi. Furthermore, collecting thousands of labeled disease images for training is both expensive and time-consuming. This paper addresses both challenges by combining neural network pruning—removing unnecessary parts of the model—with few-shot learning, which enables the model to learn from limited examples. This paper proposes Disease-Aware Channel Importance Scoring (DACIS), a method that identifies which parts of the neural network are most important for distinguishing between different plant diseases, integrated into a three-stage Prune-then-Meta-Learn-then-Prune (PMP) pipeline. Experiments on PlantVillage and PlantDoc datasets demonstrate that the proposed approach reduces model size by 78% while maintaining 92.3% of the original accuracy, with the compressed model running at 7 frames per second on a Raspberry Pi 4, making real-time field diagnosis practical for smallholder farmers.
@@ I-B Research Gap and Contributions
  • Disease-Aware Channel Importance Scoring (DACIS): A channel importance metric combining gradient sensitivity, activation variance, and Fisher’s discriminant ratio. Scope: This is an empirically-motivated heuristic combination of known metrics, not a theoretically novel scoring function. The contribution is demonstrating its effectiveness for disease classification pruning, not claiming fundamental novelty in the individual components.
  • Prune-then-Meta-Learn-then-Prune (PMP) Pipeline: A three-stage training procedure interleaving pruning with meta-learning. Scope: This is an engineering pipeline, not a theoretical framework. Results show it outperforms single-stage alternatives on the benchmark.
@@
The remainder of this paper proceeds as follows: Section 2 situates this work within the landscape of related research, identifying specific limitations that motivate the approach. Section 3 formalizes the problem setting and introduces the mathematical framework. Section 4 details the DACIS scoring mechanism and PMP training pipeline. Section 5 presents comprehensive experimental validation across multiple datasets and evaluation protocols. Section 6 discusses practical deployment considerations and limitations, and Section 7 concludes with directions for future investigation.
@@
Figure 2 presents a high-level overview of the proposed framework, illustrating the integration of disease-aware pruning with meta-learning.
@@
Figure 2: Overview of the PMP-DACIS framework.
@@ II-A Neural Network Pruning Methodologies
  • Reconstruction Error: Channels are pruned to minimize the reconstruction error of subsequent layer activations, formulated as a LASSO regression problem [8].
@@ II-C1 Architectural Efficiency in Plant Pathology
II-C2 Transfer Learning and Domain Adaptation
Transfer learning has become indispensable for plant disease detection, particularly when labeled training data is limited. Pre-training on ImageNet-scale datasets provides generalized feature representations that transfer effectively to agricultural domains. Recent studies demonstrate that integrating transfer learning with fine-tuning strategies significantly improves model robustness across diverse crop species and lighting conditions. The combination of transfer learning with SE-MobileNet architectures achieves 99.78% accuracy on curated backgrounds and 99.33% on heterogeneous backgrounds [19], showcasing the critical importance of domain adaptation mechanisms.
@@ II-D Position of This Work
TABLE I: Comparison with Representative Prior Work
@@ III-A Notation
TABLE II: Summary of Notation
@@ III-B Problem Formulation: Shot-Adaptive Model Selection
This section analyzes the relationship between data availability and optimal model capacity, which is termed Shot-Adaptive Model Selection (SAMS).
@@ III-C Hierarchical Disease Taxonomy
Figure 4: Hierarchical disease taxonomy guiding pruning protection.
Coarse level $\mathcal{V}_{1}$: Pathogen types (288 channels) receive full protection.
Medium level $\mathcal{V}_{2}$: Symptom types receive partial protection.
Fine level $\mathcal{V}_{3}$: Specific diseases are primary pruning candidates.
@@ III-D Uncertainty Quantification in Low-Data Regimes
III-E Disease-Aware Channel Importance Scoring (DACIS)
To identify and preserve the most diagnostically relevant features, the Disease-Aware Channel Importance Score (DACIS) is proposed. Unlike conventional pruning metrics that rely solely on weight magnitude or generic activation statistics, DACIS explicitly incorporates disease class separability.
@@ III-E4 Empirical Validation of Assumptions and Limitations
TABLE III: Theoretical Assumption Validation Summary
@@ III-E5 Disease Taxonomy Construction
Grad-CAM Alignment: CNN attention overlap with pathologist-annotated diagnostic regions was quantified using Intersection-over-Union (IoU). Mean IoU = 0.62 across 200 test images (3 pathologists annotating each), indicating moderate alignment. Channels with high $\mathcal{D}$ scores showed 23% higher IoU (mean = 0.76) than low-$\mathcal{D}$ channels (mean = 0.52), with $p<0.01$.
@@ III-F Layer-Adaptive Pruning Ratios
III-G The Prune-then-Meta-Learn-then-Prune (PMP) Framework
The compression strategy, the Prune-then-Meta-Learn-then-Prune (PMP) framework, is designed to resolve the conflict between pre-training objectives and few-shot adaptation needs. By interleaving pruning with meta-learning, the final compressed architecture is optimized for the specific distribution of few-shot tasks.
@@ III-G1 Theoretical Justification for Three Stages
Why not four or more stages? 4-stage (P-M-P-M) and 5-stage (P-M-P-M-P) variants were evaluated. Results in Table VII show diminishing returns: 4-stage achieves +0.3% over 3-stage while increasing training time by 45%, and 5-stage shows no improvement (+0.1%) with 78% longer training. Among configurations evaluated, three stages represent a practical trade-off balancing accuracy and computational cost. Asymmetric patterns (e.g., P-M-M-P, P-P-M-P) and continuous pruning during meta-learning were not evaluated and remain directions for future work.
@@
Algorithm 1 PMP Framework
@@ IV-A1 PlantVillage Dataset
    Severity Stratification: Classes organized by disease progression—early (0-25% affected tissue), mid (25-60%), and late (60-100%) stages. Models trained on early-stage samples are evaluated on late-stage presentations, testing symptom progression generalization.
@@ IV-A2 PlantDoc Dataset
PlantDoc [2] contains 2,598 in-the-wild images across 27 disease classes, capturing the visual complexity of field conditions. Seven classes are reserved as novel categories for few-shot evaluation.
@@ IV-B Implementation Details
  • ProtoNet: Implemented following [4] with Euclidean distance; validated against original paper’s mini-ImageNet results ($\pm$0.5% match).
@@ IV-C Evaluation Metrics
Definition 4 (Few-Shot Stability Index).
@@
where $\sigma_{\text{acc}}$ and $\mu_{\text{acc}}$ are standard deviation and mean accuracy across 1000 randomly sampled support sets. Higher FSI indicates more stable performance.
@@ IV-D Main Results
TABLE V: Few-Shot Classification Accuracy (%) on PlantDoc (In-the-Wild). Values show mean $\pm$ episode-level std. dev. across 1000 episodes.
@@
Note on Uncertainty: All accuracy values should be interpreted with $\pm 2.3\%$ episode-level uncertainty (1000-episode standard deviation), in addition to the $\pm 0.04\%$ fold-level variance reported in Table XIV. Episode-level variance reflects inherent few-shot task variability.
@@
IV-E Ablation Studies
@@
TABLE VI: Ablation Study on PlantVillage 5-Way 5-Shot. Values show mean $\pm$ episode-level std. dev.
@@
TABLE VII: Ablation: Number of Pruning Stages (30% params). Values show mean $\pm$ episode-level std. dev.
@@
IV-F Additional Ablations
@@
TABLE VIII: Additional Ablation Studies (5-Way 5-Shot, 30% params)
@@
Key findings: (1) Fisher discriminant $\mathcal{D}$ is the most critical component; removing it causes the largest drop (-4.8%). (2) Fisher outperforms alternative discriminability metrics (MMD, KL) by 1.2-2.4%. (3) The two-stage pruning schedule outperforms both gradual and one-shot alternatives. (4) Meta-learning is moderately sensitive to $\alpha$ (inner loop rate); $\alpha=0.01$ is near-optimal.
IV-G DACIS Hyperparameter Sensitivity Analysis
@@ IV-G1 Methodology
TABLE IX: DACIS Weight Sensitivity Analysis (5-Way 5-Shot)
@@ IV-G2 Key Findings
IV-H Cross-Stage Generalization Analysis
@@
[…]
Method $\rightarrowE E\rightarrowL$ CSG
@@ IV-I Stability Analysis
Figure 7: Few-Shot Stability Index comparison across methods. Higher FSI values indicate more consistent performance across different support set samplings.
@@
IV-J Multi-Resolution Robustness
@@
TABLE XI: Multi-Resolution Evaluation (Train: $\times224$)
@@
IV-K Ablation Study: Component Contribution Analysis
@@
IV-L Statistical Significance and Robustness
IV-L1 Testing Methodology
@@ IV-L2 Variance Decomposition
TABLE XIII: Statistical Significance (Paired t-tests, n=1000 episodes)
@@
IV-M Five-Fold Cross-Validation
@@
TABLE XIV: 5-Fold Cross-Validation (5-Way 5-Shot). Note: Values show fold-level variance ($\pm$0.04%). Episode-level variance is higher ($\pm$2.3%), reflecting inherent few-shot task variability.
@@
IV-N Per-Class Performance Analysis
@@
TABLE XV: Per-Class Performance (15-Way Classification)
@@
TABLE XVI: Few-Shot Classification Performance (Episodic Evaluation). $\sigma_{ep}$ denotes episode-level standard deviation across 1000 episodes.
@@
TABLE XVII: SOTA Few-Shot Methods Comparison
@@
TABLE XVIII: Fair Comparison at Equivalent Compression (30% params). Values without $\pm$ are single-run results; all methods share identical episode-level variance ($\pm$2.1-2.8% depending on shot count).
@@ IV-P Computational Efficiency
Figure 8: Layer-wise channel retention. DACIS preserves early-layer texture features (75%) while aggressively pruning semantic layers (20%) where disease-specific channels concentrate.
IV-Q Training Convergence and Learning Dynamics
@@
Figure 9: Training convergence over 25 epochs. Validation peaks at epoch 14 (99.78%) with tight post-epoch-10 convergence (¡ 0.5% variance).
IV-R Multi-Resolution Robustness Analysis

### v2

@@ Abstract
A key challenge in agricultural AI is deploying disease detection systems in remote fields with limited access to laboratories or high-performance computing (HPC) resources. While deep learning (DL) models, specifically deep convolutional networks, achieve high accuracy in identifying plant pathologies from leaf imagery, their memory footprints and computational demands limit edge deployment on devices constrained by battery life, processing power, and connectivity, such as Raspberry Pi. Few-shot learning (FSL) paradigms offer a compelling solution to the data scarcity problem inherent in agricultural applications, where obtaining labeled samples for novel disease variants proves both costly and time-sensitive. This work introduces a framework combining pruning with meta-learning for agricultural disease classification, addressing the tension between generalization capability and deployment feasibility. The proposed approach combines a novel Disease-Aware Channel Importance Scoring (DACIS) mechanism with a three-stage Prune-then-Meta-Learn-then-Prune (PMP) pipeline. Experiments on PlantVillage and PlantDoc datasets demonstrate that the proposed approach reduces model size by 78% while maintaining 92.3% of the original accuracy. The compressed model achieves 7 frames per second (FPS) on a Raspberry Pi 4, enabling practical real-time field diagnosis for smallholder farmers.
@@ I-B Research Gap and Contributions
  • Disease-Aware Channel Importance Scoring: A channel importance metric combining gradient sensitivity, activation variance, and Fisher’s discriminant ratio. Scope: This is an empirically-motivated heuristic combination of known metrics, not a theoretically novel scoring function. The contribution is demonstrating its effectiveness for disease classification pruning, not claiming fundamental novelty in the individual components.
  • Prune-then-Meta-Learn-then-Prune Pipeline: A three-stage training procedure interleaving pruning with meta-learning. Scope: This is an engineering pipeline, not a theoretical framework. Results show it outperforms single-stage alternatives on the benchmark.
@@
The remainder of this paper proceeds as follows: Section 2 situates this work within the landscape of related research, identifying specific limitations that motivate the approach. Section 3 formalizes the problem setting and introduces the mathematical framework. Section 4 details the disease-aware channel importance scoring mechanism and the Prune-then-Meta-Learn-then-Prune training pipeline. Section 5 presents comprehensive experimental validation across multiple datasets and evaluation protocols. Section 6 discusses practical deployment considerations and limitations, and Section 7 concludes with directions for future investigation.
Figure 1 illustrates representative disease samples from the PlantVillage dataset, demonstrating the visual complexity that motivates the disease-aware approach. Figure 2 presents a high-level overview of the proposed framework, illustrating the integration of disease-aware pruning with meta-learning.
@@
Figure 2: Overview of the Prune-then-Meta-Learn-then-Prune – Disease-Aware Channel Importance Scoring (PMP-DACIS) framework.
@@ II-A Neural Network Pruning Methodologies
  • Reconstruction Error: Channels are pruned to minimize the reconstruction error of subsequent layer activations, formulated as a Least Absolute Shrinkage and Selection Operator (LASSO) regression problem [8].
@@ II-C1 Architectural Efficiency in Plant Pathology
II-C2 Transfer Learning, Domain Adaptation, and Research Gaps
Transfer learning has become indispensable for plant disease detection, particularly when labeled training data is limited. Pre-training on ImageNet-scale datasets provide generalized feature representations that transfer effectively to agricultural domains. Recent studies demonstrate that integrating transfer learning with fine-tuning strategies significantly improves model robustness across diverse crop species and lighting conditions. The combination of transfer learning with SE-MobileNet architectures achieves 99.78% accuracy on curated backgrounds and 99.33% on heterogeneous backgrounds [19], showcasing the critical importance of domain adaptation mechanisms.
@@ II-D Position of This Work
TABLE I: Comparison with Representative Prior Work. Methods compared across pruning strategy, meta-learning integration, disease-awareness, edge deployment, and plant pathology inclusion. PMP-DACIS uniquely combines all five aspects.
@@ III-A Notation and Symbols
TABLE II: Summary of Notation. Mathematical symbols, sets, and operators used throughout the paper, including model parameters $\theta$, meta-learning gradients $G_{\text{meta}}$, and disease taxonomy indices.
@@ III-B Problem Formulation: Shot-Adaptive Model Selection
This section analyzes the relationship between data availability and optimal model capacity, which is termed Shot-Adaptive Model Selection.
@@ III-C Hierarchical Disease Taxonomy
Figure 4: Hierarchical Disease Taxonomy Guiding Pruning Protection (ResNet-18). Coarse: 288 channels protecting pathogen types (full protection). Medium: symptom-type features (partial protection). Fine: species-specific diseases (pruning candidates). Demonstrates disease-taxonomy alignment in channel selection.
@@ III-D Uncertainty Quantification in Low-Data Regimes
III-E DACIS
To identify and preserve the most diagnostically relevant features, DACIS is proposed. Unlike conventional pruning metrics that rely solely on weight magnitude or generic activation statistics, DACIS explicitly incorporates disease class separability.
@@ III-E4 Empirical Validation of Assumptions and Limitations
TABLE III: Theoretical Assumption Validation Summary. Quantitative verification of assumptions (A1-A4) from Section 3 using PlantVillage meta-training set. Validation status: met/partially-met/not-met with confidence metrics and supporting evidence.
@@ III-E5 Disease Taxonomy Construction
Gradient-weighted Class Activation Mapping (Grad-CAM) Alignment: CNN attention overlap with pathologist-annotated diagnostic regions was quantified using Intersection-over-Union (IoU). Mean IoU = 0.62 across 200 test images (3 pathologists annotating each), indicating moderate alignment. Channels with high $\mathcal{D}$ scores showed 23% higher IoU (mean = 0.76) than low-$\mathcal{D}$ channels (mean = 0.52), with $p<0.01$.
@@ III-F Layer-Adaptive Pruning Ratios
III-G The PMP Framework
The compression strategy, the PMP framework, is designed to resolve the conflict between pre-training objectives and few-shot adaptation needs. By interleaving pruning with meta-learning, the final compressed architecture is optimized for the specific distribution of few-shot tasks.
@@ III-G1 Theoretical Justification for Three Stages
Why not four or more stages? 4-stage (Prune (P)-then-Meta-Learn (M)-then-Prune (P)-then-Meta-Learn (M)) and 5-stage (Prune (P)-then-Meta-Learn (M)-then-Prune (P)-then-Meta-Learn (M)-then-Prune (P)) variants were evaluated. Results in Table VII show diminishing returns: 4-stage achieves +0.3% over 3-stage while increasing training time by 45%, and 5-stage shows no improvement (+0.1%) with 78% longer training. Among configurations evaluated, three stages represent a practical trade-off balancing accuracy and computational cost. Asymmetric patterns (e.g., P-M-M-P, P-P-M-P) and continuous pruning during meta-learning were not evaluated and remain directions for future work.
@@
Algorithm 1 PMP Framework: A Three-Stage Pipeline for Few-Shot Plant Disease Detection
@@ IV-A1 PlantVillage Dataset
    Severity Stratification: Classes are organized by disease progression—early (0–25% affected tissue), mid (25–60%), and late (60–100%) stages. Models trained on early-stage samples are evaluated on late-stage presentations, testing symptom progression generalization.
@@ IV-A2 PlantDoc Dataset
PlantDoc [2] contains 2,598 in-the-wild images across 27 disease classes, capturing the visual complexity of field conditions. Seven classes are reserved as novel categories for few-shot evaluation. Table V presents the few-shot classification results on this in-the-wild dataset, demonstrating the robustness of the proposed approach under realistic field conditions.
@@ IV-B Implementation Details
  • Prototypical Networks (ProtoNet): Implemented following [4] with Euclidean distance; validated against original paper’s mini-ImageNet results ($\pm$0.5% match).
@@ IV-C Evaluation Metrics
Definition 4 (Feature Stability Index).
@@
where $\sigma_{\text{acc}}$ and $\mu_{\text{acc}}$ are standard deviation and mean accuracy across 1000 randomly sampled support sets. Higher Feature Stability Index (FSI) values indicate more consistent performance.
@@ IV-D Main Results
TABLE V: Few-Shot Classification Accuracy (%) on PlantDoc (In-the-Wild). Values show mean $\pm$ episode-level std. dev. across 1000 episodes (n=1000).
@@
Note on Uncertainty: All accuracy values should be interpreted with $\pm 2.3$% episode-level uncertainty (1000-episode standard deviation), in addition to the $\pm 0.04$% fold-level variance reported in Table XIV. Episode-level variance reflects inherent few-shot task variability.
@@
IV-E Comprehensive Ablation Analysis: Component and Stage Contributions
@@
TABLE VI: Ablation Study on PlantVillage 5-Way 5-Shot. Values show mean $\pm$ episode-level std. dev. on pruned ResNet-18 (30% compression) with DACIS component contributions.
@@
TABLE VII: Ablation: Number of Pruning Stages (30% params, PlantVillage 5-Way 5-Shot). Compares 1-stage (pre-only), 2-stage (pre+post), and 3-stage (pre+meta+post) PMP framework. Values show mean $\pm$ episode-level std. dev.
@@
IV-E1 Alternative Metrics and Component Combinations
@@
TABLE VIII: Additional Ablation Studies (5-Way 5-Shot, 30% params). Components: $\mathcal{G}$ = gradient norm, $\mathcal{V}$ = variance, $\mathcal{D}$ = Fisher discriminant. All evaluated on PlantVillage with ResNet-18 backbone.
@@
Key findings: (1) Fisher discriminant $\mathcal{D}$ is the most critical component; removing it causes the largest drop (-4.8%). (2) Fisher outperforms alternative discriminability metrics (Maximum Mean Discrepancy, Kullback-Leibler divergence) by 1.2-2.4%. (3) The two-stage pruning schedule outperforms both gradual and one-shot alternatives. (4) Meta-learning is moderately sensitive to $\alpha$ (inner loop rate); $\alpha=0.01$ is near-optimal.
IV-F DACIS Hyperparameter Sensitivity Analysis
@@ IV-F1 Hyperparameter Tuning Methodology
TABLE IX: DACIS Weight Sensitivity Analysis (5-Way 5-Shot) with Selected Weights $\lambda_{1}=0.3,\lambda_{2}=0.2,\lambda_{3}=0.5$
@@ IV-F2 Key Findings
IV-G Cross-Stage Generalization Analysis
@@
TABLE X: Cross-Stage Generalization: Early→Late Severity Transfer (5-Way 5-Shot). Measures model capability to generalize from mild to severe disease symptoms across plant progression stages, evaluating disease taxonomy-guided pruning effectiveness.
Method $\rightarrowE E\rightarrowL$ Cross-Stage Gen.
@@ IV-H Stability Analysis
Figure 7: Few-Shot Stability Index (FSI) Comparison on PlantVillage. Higher FSI values indicate more consistent performance across different support set samplings (20 trials per method). PMP-DACIS achieves highest FSI = 0.92 vs ProtoNet (0.71), validating disease-aware feature preservation.
@@
IV-H1 Multi-Resolution and Robustness Analysis
@@
TABLE XI: Multi-Resolution Evaluation (5-Way 5-Shot). Model trained at 224×224; tested across 64, 112, 224, 256, 448 pixel resolutions. Measures generalization robustness on PlantVillage dataset.
@@
IV-H2 Component Contribution Analysis
@@
IV-I Statistical Significance and Robustness
IV-I1 Testing Methodology
@@ IV-I2 Variance Decomposition
TABLE XIII: Statistical Significance (Paired t-tests, n=1000 episodes) with Bonferroni and Holm-Bonferroni corrections applied (adjusted threshold $\alpha=0.00037$)
@@
IV-I3 Five-Fold Cross-Validation
@@
TABLE XIV: 5-Fold Cross-Validation (5-Way 5-Shot on PlantVillage). Fold-level variance ($\pm$0.04%) is negligible; episode-level variance ($\pm$2.3%) reflects inherent few-shot task sampling variability. Validates robustness across data splits.
@@
IV-J Per-Class Performance Analysis
@@
TABLE XV: Per-Class Performance (15-Way Classification) on PlantVillage with DACIS-pruned ResNet-18. Precision, Recall, F1-score for each of 15 disease/health classes.
@@
TABLE XVI: Few-Shot Classification Performance (Episodic Evaluation on PlantVillage). $\sigma_{ep}$ denotes episode-level standard deviation across 1000 episodes. 5-Way k-Shot evaluation with varying support set sizes.
@@
TABLE XVII: SOTA Few-Shot Methods Comparison on PlantVillage 5-Way setting. Params(M) = parameter count in millions; fair comparison at equivalent model capacity (0.11M baseline). PMP-DACIS achieves 83.2% at 30% compression.
@@
TABLE XVIII: Fair Comparison at Equivalent Compression (30% params). All methods use ResNet-18 backbone pruned to 3.36M parameters. Values without $\pm$ are single-run results; episode-level variance is $\pm$2.1-2.8% (controlled across methods).
@@ IV-L Computational Efficiency
Figure 8: Layer-wise Channel Retention (30% compression). DACIS preserves early-layer texture features (75% retention) while aggressively pruning semantic layers (20% retention) where disease-specific channels concentrate. Data from ResNet-18 layer-wise analysis.
IV-L1 Training Convergence and Learning Dynamics
@@
Figure 9: Training Convergence over 25 Epochs on Meta-Training Set. Validation peaks at epoch 14 (99.78%) with tight post-epoch-10 convergence (<<0.5% variance). Indicates stable optimization without overfitting, validating early stopping criterion.
IV-L2 Multi-Resolution and Resolution Robustness Analysis